# Lemmanaid: Neuro-Symbolic Lemma Conjecturing

### v3

@@ 3.2 Template Prediction via Lemmanaid’s Neural Engine
3.3 Direct Neural Lemma Conjecturing
While Lemmanaid employs a neural engine to generate templates and a symbolic engine to instantiate those templates into lemmas, there is another approach one can take, which is to use a neural engine to directly generate lemmas. To create such a baseline, we adapt Lemmanaid’s neural engine, by instead fine-tuning it on tuples of the form (symbols, context, lemma) instead of a template. Revisiting the running example again, the lemma octo_distrib_left now results in a datapoint of the shape:
Input: [Symbols: *, +] [Types of *, +] [Defs of *, +] ; Output: [$(a*(b+c)=a*b+a*c$]
@@ 3.4 Dataset
We train and evaluate Lemmanaid on mathematical libraries from the Isabelle proof assistant. First, the Isabelle/HOL library, which contains formalizations based on higher order logic for a range of mathematics (e.g. number theory, analysis, algebra, set theory). Second, we also use the Isabelle Archive of Formal Proofs (AFP), which is a large collection of formalizations from the research community including mathematics, computer science and logic. At the time of writing, the AFP includes about 285,200 lemmas in close to 900 theories. We extract templates for the lemmas in the above libraries as described in §3.1. In total, this results in a dataset of 62,816 data-points from the HOL library and 206,304 data-points from the AFP [^5] Some theories could not be processed by the batch method we used, due to technical issues with theory imports. Hence, our dataset does not cover everything from the AFP..
@@ 1. RQ1
    How does our neuro-symbolic approach, Lemmanaid, compare with neural approaches and existing symbolic approaches? How does different contextual information (such as types and definitions) impact performance?
@@ 2. RQ2
    How does an LLM compare to Lemmanaid’s symbolic engine in its ability to fill in lemma templates?
@@ 3. RQ3
    How well does an LLM understand the lemma template language?
  4. RQ4
    What kinds of lemmas is Lemmanaid able to generate?
@@ 4. RQ4
Table 1: Percentage of gold-standard lemmas discovered for Lemmanaid and a neural only version. We include results for the symbolic QuickSpec tool only on Octonions which is in its scope. Results are for 5 predictions (greedy decoding & beam search with beam size 4).
@@
Lemmanaid (types + defs) 38.8% 25.4% 53.7% 38.6% 29.1% 47.4%
Lemmanaid (types) 35.6% 25.4% 58.5% 39.5% 26.5% 52.0%
Lemmanaid (defs) 33.0% 13.8% 45.7% 33.9% 14.7% 41.4%
Neural (types + defs) 30.7% 13.8% 34.6% 29.1% 14.9% 39.1%
Neural (types) 31.5% 14.4% 43.4% 28.9% 14.6% 41.1%
Neural (defs) 26.7% 8.6% 30.6% 28.5% 10.4% 36.3%
Lemmanaid Combined 46.5% 30.9% 68.0% 48.2% 34.6% 65.7%
Neural Combined 39.7% 19.5% 49.1% 38.9% 20.8% 56.6%
Combined 50.7% 32.9% 71.4% 52.0% 37.4% 74.0%
@@
4.1 Experimental Setup
Baselines
While prior work (Gauthier & Urban, 2023, 2025) has explored direct neural lemma conjecturing, there are no existing neural-based tools or models for Isabelle to compare against. We train our own neural baselines (Section 3.3), and go beyond prior work by predicting the lemma objects. Models trained to predict the lemma object are denoted by “neural” in tables in the evaluation. The state-of-the-art symbolic tool for lemma conjecturing is QuickSpec (Smallbone et al., 2017). QuickSpec is limited in that it can only generate equational lemmas as its conjecturing algorithm is based around enumerative synthesis of terms and the construction of equivalence classes via automated testing.
Benchmarks
We derive our datasets from the Isabelle/HOL library and the Archive of Formal Proofs (AFP) (recall §3.4) and create multiple train/validation/test sets from these libraries. We create a file-wise split of the HOL library so that we may evaluate the in-distribution capabilities of LLM-based approaches for lemma conjecturing tasks. The training, validation and test sets are called HOL-train (57,576 datapoints), HOL-val (500 datapoints), and HOL-test (4,740 datapoints), respectively.
Next, we supplement HOL-train and HOL-val with all projects from the AFP2024 that are published prior to 2024 to create HOL+AFP-train and HOL+AFP-val. We then create a new test set called AFP-test comprised of 27 AFP projects published in 2024 (and thus disjoint from those in HOL+AFP-train and HOL+AFP-val). More information about results on the different projects in AFP-test can be found in A.3. Training models on HOL+AFP-train and evaluating on HOL-test allows us to understand the effect of more training data as opposed to training on only HOL-train. On the other hand, training models on HOL+AFP-train and evaluating on AFP-test allows us to evaluate an out-of-distribution task. Importantly, evaluating on AFP-test allows us to mitigate the risk of test leakage as the models we use have publicly reported pretraining cutoff dates in 2023. AFP-test consists of 16,362 datapoints. We separately test on the Octonions project (Koutsoukou-Argyraki, 2018), which contains 350 lemmas, from the AFP and leave it out of all training sets in order to compare Lemmanaid, direct neural lemma conjecturing models, and symbolic tool, QuickSpec. We choose the Octonion theory for this comparison as it consists of equational lemmas, which is the domain QuickSpec supports. Recall that QuickSpec can only conjecture equational lemmas on computable functions which limits its applicability to a smaller subset of Isabelle theories.
We enumerate the lemmas appearing in a given project by compiling and processing a theory and counting the number of lemma objects that exist. We do this by using the Find_Theorems tool in Isabelle which retrieves theorems from a proof context, and we only keep the theorems defined in the active theory. The implications of this enumeration are discussed in A.2.
Metrics
We define multiple metrics used throughout our evaluation. In our approach, we have a set of lemma-prediction tasks. In each such task, there is one gold-standard lemma, and the method being assessed generates a set of predicted lemmas. We define lemma success rate as the percentage of these lemma prediction tasks for which the given method is able to successfully generate (as part of the set of lemmas it generates) the gold-standard lemma (where we compare lemmas syntactically). This overall metric measures the performance of a method end-to-end. For Lemmanaid’s neural engine, we want to measure template success rate, which is the percentage of template-prediction tasks for which it predicts the correct gold-standard template. For the symbolic engine, given a template, the symbolic engine generates many instantiations, and we consider the instantiation task to be successful if one of the generated instantiations matches the gold-standard lemma (syntactically). The instantiation rate measures the percentage of instantiation tasks that the symbolic engine can perform successfully. We also use this metric in RQ2 to measure an LLM’s ability to instantiate templates. To answer RQ3, we need to measure how well LLMs are able to generate templates from lemma objects. In this case, a task consists of predicting a template from a lemma object. We define abstraction rate as the percentage of such lemma-template-from-object prediction tasks the LLM is able to perform.
Models and Inference
For all tasks, we use deepseek-coder-1.3b-base as the pretrained model that we fine-tune. In early experiments, we explored the use of Llama-3.2-1b, but found that all methods performed worse. One benefit of using smaller models in Lemmanaid is that it supports a more realistic setup for actual users wanting to use a conjecturing tool and run it locally on their machine. See A.4 for some details about computing resources and replication parameters used in our evaluation. At inference time, we use greedy decoding to obtain a template prediction from Lemmanaid’s LLM. We also use beam search with beam size equal to 4. For our RQ1 evaluation, we use both decoding strategies for all neural models, including neural baselines, and thus all methods with neural models have the same LLM inference budget of 5. For all other RQs, greedy decoding is used for LLM inference.
Checking Outputs
To check the correctness of a predicted template, we use the exact match (string similarity) of the predicted template and the abstracted template of the gold-standard lemma we want to recover. This is sufficient for templates given the way in which we define the template language. Exact match is not sufficient for comparing lemma objects because it requires identical variable names and therefore does not account for alpha equivalence of the terms. We want to count a generated lemma as matching the gold-standard even if it uses different variable names. We parse the gold-standard lemma object and the predicted lemma object in the context of a given theory, as some symbols are only defined in that context. We perform term comparison, where we traverse the tree, checking for equivalences on left and right, respectively. Since alpha renaming is already implemented in Isabelle, we use this for variable comparison. Note that this cannot be done easily outside of Isabelle because we must know a given variable’s scope. Since instantiation of a predicted template via Lemmanaid’s symbolic engine produces multiple candidate lemmas, we iterate over them and check each against the gold-standard lemma. Lemmanaid’s symbolic engine’s instantiation timeout is set to 60s.
4.2 RQ1: Neuro-symbolic vs Neural vs Symbolic
@@
To evaluate Lemmanaid and compare it against neural-only baselines and symbolic methods, we train Lemmanaid’s neural model and the baseline models on either HOL-train or HOL+AFP-train, obtain multiple variants, and evaluate them across different test sets (Table 1). For all models, we compare the results obtained when various contextual information is included in the input: definitions, type information, or both (as described in 3.2). For each model, we get 5 predictions: 1 using greedy decoding, and 4 using beam search (with beam size of 4). We break down the results for each decoding method in Appendix A.5.
@@
We see that Lemmanaid outperforms the respective neural baselines on all test sets. We also see that neural methods are somewhat complementary to Lemmanaid, taken as an ensemble they discover even more lemmas. The inclusion of type information is greatly beneficial to both Lemmanaid and the neural baseline method. We see that in some cases the success rate is higher when only type information is included and definitions are excluded, while in others including both is beneficial, although differences are relatively small. We see that on AFP-test, the performance drops for all variants compared to their results on HOL-test. This is unsurprising, as the lemmas in AFP projects are more diverse than those in HOL. Also, the input to the LLMs may not include enough contextual information as retrieving definitions in AFP theories requires retrieving across dependencies, though we always account for HOL as a dependency. We see that when trained on HOL+AFP-train, performance improves on AFP-test, in particular for Lemmanaid where we see an increase from 25.4% to 29.1%, while the neural baseline only increases from 14.4% to 14.9%.
For Octonions, Lemmanaid models outperform their respective neural baselines while both greatly outperform QuickSpec. Similar to other datasets, an ensemble of Lemmanaid and neural models provides a significant improvement, with overall lemma success rate up to 74%. However, QuickSpec is complementary to both models trained on HOL-train and HOL+AFP-train, improving overall lemma success rate to 74.6 % and 76.9% respectively.
[…]
4.3 RQ2: Template instantiation by Lemmanaid’s symbolic engine vs an LLM
As discussed in Section 3.1, Lemmanaid instantiates its predicted templates symbolically. We also noted some limitations of this instantiation and how it is subject to timeouts. Alternatively, we can train an LLM to instantiate templates. Table 2 displays the template instantiation rate for Lemmanaid’s symbolic engine and an LLM trained to instantiate templates when both are given the gold-standard lemma template and the symbols appearing in the gold-standard lemma. We observe that Lemmanaid’s symbolic engine has a much greater instantiation rate (89.1%) than that of an LLM (66.9%). However, interestingly, the LLM recovers some lemmas that we symbolically fail to recover, so when we consider lemmas recovered either symbolically or neurally, the rate is 92.2%.
Method Instantiation Rate
LLM 66.9%
Lemmanaid 89.1%
Combined 92.2%
4.4 RQ3: Template abstraction performance of LLM
Table 3: Template abstraction rate for LLMs (fine-tuned on HOL-train) with different inputs evaluated on HOL-test.
LLM Input Abstraction Rate
Lemma Object 88.9%
+ Symbols + Types 92.0%
+ Symbols + Types + Defs 90.3%
Combined 94.5%
Table 4: Statistics on HOL-test, AFP-test, and Octonions, and the successes of Lemmanaid and neural baselines on each dataset. Lemmanaid and neural models here are trained on HOL-train and use greedy decoding.
Lemma Count Equational Template Length Lemma Length # Symbols
Min 25% Median 75% Max Min 25% Median 75% Max Mean Max
HOL-test 4740 2242 9 49 82 127 9014 7 57 91 150 10454 4.2 100
Lemmanaid Success 1323 734 11 33 53 77 289 10 43 65 90 325 3.0 13
Neural Success 1135 629 11 37 55 77 289 11 45 65 92 347 3.1 15
AFP-test 16362 10582 7 37 72 142 10794 7 56 88 152 7530 4.2 31
Lemmanaid Success 2420 1894 7 23 29 43 210 10 37 52 72 246 2.7 10
Neural Success 1594 1142 11 23 33 45 173 12 34 48 68 192 2.7 13
Octonions 350 262 17 29 49 75 688 11 33 55 79 391 3.8 21
Lemmanaid Success 143 107 17 23 39 59 109 11 25 43 60 133 2.5 6
Neural Success 106 66 19 19 44 67 109 15 24 47 73 133 2.4 7
To assess how well LLMs understand the language of lemma templates, we train an LLM to abstract templates from lemma objects. We have three variants of this task: (1) only lemma objects as input, (2) lemma objects, symbols, and their types as inputs, and (3) which is similar to (2) with the addition of definitions (Table 3). Models perform similarly, and there are some non-overlapping examples that each model uniquely predicts so the combination has 93.7% abstraction rate. Notably, a point of confusion for models is operators with custom syntax. The use of custom syntax introduces ambiguity about how the arguments of a function are ordered. Custom syntax can possibly make some arguments implicit, which introduces further ambiguity. Overall, these results indicate that LLMs can understand the template language well.
4.5 RQ4: Qualitative analysis
To assess the utility of our tools, we study the characteristics of HOL-test, Octonions and AFP-test, as well as the subsets of lemmas that Lemmanaid and neural baselines successfully generate. For each subset, we consider the number of equational lemmas, the length (in characters) of the templates and lemmas, and the number of symbols appearing in the lemmas (Table 4). Both Lemmanaid and the neural baselines demonstrate an ability to predict equational and non-equational lemmas. They successfully discover lemmas and templates of different lengths but show greater ability to handle shorter ones. This is to be expected because some contextual information for longer lemmas would be cutoff due to our max sequence length (where we limit the number of input and output tokens) and they are likely more complex for an LLM to reason about. See A.7 for some examples that Lemmanaid synthesizes that other methods do not. See A.8 for some examples of how Lemmanaid succeeds and fails.

### v5

@@ 3.2 Lemmanaid’s Neural Engine
3.3 Direct Neural Conjecturing
While Lemmanaid employs a neural engine to generate templates and a symbolic engine to instantiate those templates into lemmas, there is another approach one can take, which is to use a neural engine to directly generate lemmas. To create such a baseline, we adapt Lemmanaid’s neural engine, by instead fine-tuning it on tuples of the form (symbols, context, lemma) instead of a template. Revisiting the running example again, the lemma octo_distrib_left now results in a data-point of the shape:
Input: [Symbols: *, +] [Types of *, +] [Defs of *, +] ; Output: [(a(b+c)=ab+ac(a*(b+c)=a*b+a*c]
@@ 3.4 Dataset
We train and evaluate Lemmanaid on mathematical libraries from the Isabelle proof assistant. First, the Isabelle/HOL library, which contains formalizations based on higher order logic for a range of mathematics (e.g. number theory, analysis, algebra, set theory). Second, we also use the Isabelle Archive of Formal Proofs (AFP), which is a large collection of formalizations from the research community including mathematics, computer science and logic. At the time of writing, the AFP includes about 285,200 lemmas in close to 900 theories. We extract templates for the lemmas in the above libraries as described in §3.1. In total, this results in a dataset of 62,816 data-points from the HOL library and 210,898 data-points from the AFP [^5] Some theories could not be processed by the batch method we used, due to technical issues with theory imports. Hence, our dataset does not cover everything from the AFP..
@@ 1. RQ1
    How does our neuro-symbolic approach, Lemmanaid, compare with neural approaches and existing symbolic approaches?
@@ 2. RQ2
    How does different contextual information (such as types and definitions) impact performance?
@@ 3. RQ3
    What characterizes the conjectures Lemmanaid is able to generate (how many can be proved/have counter examples)?
4.1 Experimental Setup
Baselines
There are no existing neural-based tools or models for Isabelle to compare against so we train our own neural baselines (Section 3.3), to suggest lemma objects. Models trained to predict the lemma object are denoted by “neural” in tables in the evaluation. The state-of-the-art symbolic tool for lemma conjecturing is QuickSpec (Smallbone et al., 2017). QuickSpec is limited in that it can only generate equational lemmas as its conjecturing algorithm is based around enumerative synthesis of terms and the construction of equivalence classes via automated testing.
Benchmarks
We derive our datasets from the Isabelle/HOL library and the Archive of Formal Proofs (recall §3.4) and create multiple train/validation/test sets from these libraries. We create a file-wise split of the HOL library so that we may evaluate the in-distribution capabilities of LLM-based approaches for lemma conjecturing tasks. The training, validation and test sets are called HOL-train (57,576 datapoints), HOL-val (500 datapoints), and HOL-test (4,740 datapoints), respectively.
Next, we supplement HOL-train and HOL-val with all projects from the AFP2024 that are published prior to 2024 to create HOL+AFP-train and HOL+AFP-val. We then create a new test set called AFP-test comprised of 27 AFP projects published in 2024 (and thus disjoint from those in HOL+AFP-train and HOL+AFP-val). More information about results on the different projects in AFP-test can be found in E.2. Training models on HOL+AFP-train and evaluating on HOL-test allows us to understand the effect of more training data as opposed to training on only HOL-train. On the other hand, training models on HOL+AFP-train and evaluating on AFP-test allows us to evaluate an out-of-distribution task (as AFP projects tend to differ greatly in topic, content, and style). Importantly, evaluating on AFP-test allows us to mitigate the risk of test leakage as the models we use have publicly reported pretraining cutoff dates in 2023, and all projects in AFP-test are published in 2024 and after. AFP-test consists of 18,975 datapoints. Note that risk for test-data leakage is bigger for the purely neural baseline, as Lemmanaid’s neural model produces our custom template format which is not explicitly present in the original data source.
We separately test on the Octonions project (Koutsoukou-Argyraki, 2018) from the AFP, which contains 350 lemmas, and leave it out of all training sets in order to compare Lemmanaid, direct neural lemma conjecturing models, and symbolic tool, QuickSpec. We choose the Octonion theory for this comparison as it consists of equational lemmas, which is the domain QuickSpec supports. Recall that QuickSpec can only conjecture equational lemmas on computable functions which limits its applicability to a smaller subset of Isabelle theories.
We enumerate the lemmas appearing in a given project by compiling and processing a theory and counting the number of lemma objects that exist. We do this by using the Find_Theorems tool in Isabelle which retrieves theorems from a proof context, and we only keep the theorems defined in the active theory.
Metrics
In each evaluation task, there is one gold-standard lemma, and the method being assessed generates a set of suggested lemmas. We define lemma success rate as the percentage of these tasks for which the given method is able to successfully generate (as part of the set of lemmas it generates) the gold-standard lemma (where we compare lemmas syntactically). This overall metric measures the performance of a method end-to-end. Note that the purely neural model and Lemmanaid are given the same neural computational budget per task. However, Lemmanaid’s symbolic engine may generate several instantiations from the same template (this is however computationally cheap in comparison to the neural inference), and we consider the instantiation task to be successful if one of the generated instantiations matches the gold-standard lemma (syntactically).
Models and Inference
We evaluate three instances of Lemmanaid using DeepSeek-coder-1.3B, DeepSeek-coder-6.7B and Llama-3.2-1B as the pretrained model that we fine-tune. These models are selected as representatives of open-weight base models performant on code-related tasks. One benefit of using smaller models in Lemmanaid is that it supports a more realistic setup for actual users wanting to use a conjecturing tool and run it locally on their machine. See Appendix B for some details about computing resources and replication parameters used in our evaluation. At inference time, we use greedy decoding to obtain a template prediction from Lemmanaid’s LLM. We also use beam search with beam size equal to 4. For our RQ1 evaluation, we use both decoding strategies for all neural models, including neural baselines, and thus all methods with neural models have the same LLM inference budget of 5.
Checking Outputs
To check the correctness of a suggested template, we use the exact match (string similarity) of the predicted template and the abstracted template of the gold-standard lemma we want to recover. This is sufficient for templates given the way in which we define the template language. Exact match is not sufficient for comparing lemma objects because it requires identical variable names and therefore does not account for alpha equivalence of the terms. We want to count a generated lemma as matching the gold-standard even if it uses different variable names. We parse the gold-standard lemma object and the predicted lemma object in the context of a given theory, as some symbols are only defined in that context. We perform term comparison, where we traverse the tree, checking for equivalences on left and right, respectively. Since alpha renaming is already implemented in Isabelle, we use this for variable comparison. Note that this cannot be done easily outside of Isabelle because we must know a given variable’s scope. Since instantiation of a predicted template via Lemmanaid’s symbolic engine produces multiple candidate lemmas, we iterate over them and check each against the gold-standard lemma. Lemmanaid’s symbolic engine’s instantiation timeout is set to 60s.
4.2 RQ1–2: Discovery of Gold Standard Lemmas
@@
Table 1: Percentage of gold-standard lemmas discovered for Lemmanaid and a neural only version. We include results for the symbolic QuickSpec tool only on Octonions which is in its scope. Results are for 5 predictions (greedy decoding & beam search with beam size 4). Note that due to limited computational resources and poor results using DeepSeek 1.3B, we did not further explore the prompting technique with only definitions for the other models.
@@
DeepSeek-coder-1.3B
Lemmanaid (types + defs) 39.3% 19.8% 57.4% 41.7% 27.7% 66.9%
Lemmanaid (types) 37.4% 19.4% 58.3% 35.5% 27.1% 55.7%
Lemmanaid (defs) 33.0% 7.1% 54.0% 35.4% 12.8% 61.4%
Neural (types + defs) 29.3% 9.6% 33.7% 30.0% 13.5% 44.0%
Neural (types) 32.3% 10.3% 42.9% 29.0% 12.6% 47.1%
Neural (defs) 29.5% 5.4% 41.4% 25.4% 6.9% 44.0%
Lemmanaid Combined 47.6% 25.9% 67.7% 49.6% 35.3% 75.7%
Neural Combined 41.1% 14.8% 53.1% 38.3% 18.6% 58.0%
Combined 52.6% 27.7% 70.9% 52.7% 37.2% 79.1%
DeepSeek-coder-6.7B
Lemmanaid (types + defs) 46.1% 25.9% 68.0% 50.0% 27.4% 76.6%
Lemmanaid (types) 43.0% 22.1% 66.9% 47.5% 28.2% 79.1%
Neural (types + defs) 40.9% 16.8% 52.6% 28.9% 20.0% 61.7%
Neural (types) 37.3% 14.7% 51.4% 36.8% 19.0% 52.9%
Lemmanaid Combined 51.7% 28.9% 76.0% 54.9% 34.2% 83.1%
Neural Combined 45.3% 20.1% 58.9% 42.5% 25.3% 65.1%
Combined 56.4% 31.2% 79.4% 58.0% 39.1% 84.6%
Llama-3.2-1B
Lemmanaid (types+defs) 31.5% 18.0% 57.4% 28.5% 21.5% 46.9%
Lemmanaid (types) 33.9% 18.6% 59.7% 34.7% 23.4% 60.6%
Neural (types+defs) 22.1% 7.3% 33.1% 18.4% 10.1% 54.6%
Neural (types) 22.2% 8.0% 41.4% 16.2% 9.8% 57.4%
Lemmanaid Combined 39.1% 20.7% 62.6% 38.3% 29.9% 65.4%
Neural Combined 28.0% 8.7% 46.9% 23.1% 12.3% 38.3%
Combined 42.4% 21.8% 63.4% 40.8% 31.5% 66.6%
@@
To evaluate Lemmanaid and compare it against neural-only baselines and symbolic methods, we train Lemmanaid’s neural model and the baseline models on either HOL-train or HOL+AFP-train, obtain multiple variants, and evaluate them across different test sets (Table 1). For each model, we compare the results obtained when various contextual information is included in the input: definitions, type information, or both (as described in 3.2). Note that adding only the definitions, but not types, led to worse performance in initial experiments with DeepSeek 1.3B, so was not attempted for the other LLMs. For each model, we get 5 predictions: 1 using greedy decoding, and 4 using beam search (with beam size of 4). We break down the results for each decoding method in Appendix D.
@@
We see that Lemmanaid outperforms the respective neural baselines on all test sets, with the best model (using DeepSeek 6.7B and HOL+AFP training) reaching 50% on HOL-test and 28.2% on AFP-test. For the smaller 1B models, the DeepSeek backend appears to consistently outperform Llama-3.2. We also see that neural methods are somewhat complementary to Lemmanaid: taken as an ensemble they discover even more lemmas, 58% on HOL-test and 39.1% on AFP-test (using DeepSeek 6.7B and HOL+AFP training). The inclusion of type information is greatly beneficial to both Lemmanaid and the neural baseline method. We see that in some cases the success rate is higher when only type information is included and definitions are excluded, while in others including both is beneficial, although the differences are relatively small. We see that on AFP-test, the performance drops for all variants compared to their results on HOL-test. This is unsurprising, as the lemmas in AFP projects are more diverse than those in HOL. Also, the input to the LLMs may not always include enough contextual information for retrieving definitions in AFP theories requiring long-distance dependencies, though we always account for HOL as a dependency.
For Octonions, Lemmanaid models (best reaching 79.1%) outperform their respective neural baselines (61.7%) while both greatly outperform QuickSpec (reaching just 22.8%). Similar to other datasets, an ensemble of Lemmanaid and neural models provides a significant improvement, with overall lemma success rate up to 84.6%. Not shown in Table 1 is that QuickSpec generates several thousand lemmas, giving it an extremely poor precision of less than 1%. This is because it has limited heuristics for judging which lemmas are interesting, and only skips lemmas that are logical consequences of existing lemmas. It also skips several useful and simple lemmas, such as $\mathit{inner}\,e_{1}\,x=\mathit{Im}_{1}\,x$, as they are considered trivial consequences of other lemmas. This is further discussed in Appendix H.
4.3 RQ3: Analysis of Generated Conjectures
In §4.2, we counted the number of gold standard lemmas discovered. Here, we break down these results further.
Applicability to Realistic Formalizations
Appendix E shows a breakdown per projects/theories in AFP-test and HOL-test, with the number of test lemmas from each and the success rate of the best-performing Lemmanaid and neural ensembles with DeepSeek 6.7B. We note that Lemmanaid performs better on almost every test theory in the test sets. We also list the topic of the different AFP formalizations in our test set, showing that they range over a variety of different topics from Computer Science, Mathematics and Logic. The results confirm that Lemmanaid’s approach to conjecturing by analogy successfully produces interesting lemmas across a wide range of real world formalizations.
Generated Conjectures
We take a closer look at the full set of conjectures generated by the best-performing instance of Lemmanaid (DeepSeek 6.7B and HOL+AFP training), under greedy decoding. We separate the generated conjectures into three main categories (true, false and open) using the Sledgehammer tool for automated proofs Blanchette et al. (2011), and counter-example finding tools Nitpick Blanchette and Nipkow (2010) and Quickcheck Bulwahn (2012). True conjectures consist of 1) the gold-standard lemmas, and 2) other conjectures proved by Sledgehammer. False conjectures are those for which we can find a counter-examples. Open conjectures are those that can neither be proved by Sledgehammer, nor have a counter-example.
For HOL-test (containing 4,740 gold-standard lemmas), Lemmanaid generates 10,314 conjectures, however 5040 of these comes from a degenerate case [^6] These come from a case of a single template consisting of 7 OR statements resulting in 5040 equivalent conjectures. . Removing these, applying deduplication (removing 164 duplicates) leaves 5,110 conjectures for which Table 2 shows a breakdown: 1043 are novel true statements which don’t exist in the target theories, and 1256 are open conjectures which require non-trivial reasoning. Appendix F contains some examples of these, and Appendix G additional statistics.
Table 2: Generated conjectures for HOL-test.
Category Size
Gold-standard 1721 (33.7%)
Other True Statements 1043 (20.4%)
False Conjectures 1090 (21.3%)
Open Conjectures 1256 (24.6%)
Total 5110